INAF queue observing September 2013

Observer: E. Sani, F. Cusano
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson
Instrument: LUCI

September 24-25

MODS1 is currently not available due to problems with the red CCD electronics. The MODS2 head electronics box is being sent to LBTO from Columbus.
At the sunset pho sky with no wind and 41% humidity.
At the beginning of night time operations the seeing is better than 0.5 on the GCS.

Project_RUN Objects Filter ExpTime Notes
Cresci_8B DSF2237b K, 210_zJHK 3h40min within the specs
Cresci_8B HIP116324 K, 210_zJHK    
Galametz_31_A1 CANDELS_UDS18 zJspec, 200_H+K 1h35min within the specs
Galametz_31_A1 HIP11809 zJspec, 200_H+K    

September 25-26

MODS1 is currently not available due to problems with the red CCD electronics.
At sunset we have to remain closed due to strong wind.
At the beginning of the observations the sky is clear.
At 5.00 UT we need to close the dome due to strong wind. When the TO tries to open again, the left door turns to be stuck. We need to call J. Urban.
The left door problem was solved, but wind gusts increased again. The TO closed the dome.

Project_RUN Objects Filter ExpTime Notes
Cresci_8B DSF2237b K, 210_zJHK 20 min within the specs
Cresci_8B Hip104320 K, 210_zJHK    

September 26-27

MODS1 is currently not available due to problems with the red CCD electronics.
At sunset we have to remain closed due to strong wind and the sky is partly cloudy.
We stay closed all the night long due to strong wind and high humidity.

September 27-28

MODS1 is currently available in red or blue grating only.
At the sunset partially cloudy and temperature close to due point.
When the TO opens the dome we ave clear sky and low wind.
At the beginning of scientific integtrartion the average seeing on the GCS is 1.7" with 90% humidity.
During the night we had 2 issue tracks, one for stuck rotaror and the over to shrink condensation from LUCI.

Project_RUN Objects Filter ExpTime Notes
Cresci_8B Hip107734 210_zJHK, H, K   bad seeing conditions K CW=2.17
Cresci_8B DSF2237b 210_zJHK, K 1h bad seeing at the beginning. Totatal oh 5h of integration reached, but some frames to be repeated. K CW=2.17
Cresci_8B DSF2237b 210_zJHK, H 2h variable seeing, but mostly within or close to the required 1"
Cresci_8B Hip18874 210_ZJHK, H, K   K CW=2.17
Galametz_31A Hip11809 200_H+K, zJspec   condensation. Loose spectra for slits from #11 to #32
Galametz_31A CANDELS_UDS18 200_H+K, zJspec 10m condensation. Loose spectra for slits from #11 to #32

September 28-29

MODS1 is currently available in red or blue grating only.
At the sunset clear sky.
At the beginning of sceintific operations the average seeing on the GCS is 0.6".
Hanging of the Telescope Control GUI. We call the SA. We figure out by specifiyng 'wait for collimation' 'OFF' in the Telescope Specific.
Condensation on the camera. We call J. Morris that have to switch on compressors. These were non-active due to the fire alarm happened in the evening.

Project_RUN Objects Filter ExpTime Notes
Cresci_8B Hip107734 210_zJHK, H, K   for DSF2237b
Cresci_8B DSF2237b 210_zJHK, K 30m within the specs
Cresci_8B DSF2237b 210_zJHK, H 3h within the specs. target Cresci_8B COMPLETED
Cresci_8A Hip107734 210_zJHK, K   for DSF2237a
Cresci_8A DSF2237a 210_zJHK, K 40m within the specs
Galametz_31_A1 Hip11809 200_H+K, zJspec   for CANDELS_USD18_1
Galametz_31_A1 CANDELS_USD18_1 200_H+K, zJspec 50m within the specs. target Galametz_31_A1 COMPLETED
Galametz_31_A2 CANDELS_USD18_2 200_H+K, zJspec 30m within the specs.
Galametz_31_A1 Hip11809 200_H+K, zJspec   for CANDELS_USD18_2

-- EleanoraSani - 29 Sep 2013
Topic revision: r2 - 02 Feb 2014, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback